OpenStudy (sawq):

Will fan and medal!!! 4w − 7k = 28 solve for k I know I'm supposed to reorder the terms and them do the opposite operation but I keep getting confused which numbers to do HELP PLS

OpenStudy (anonymous):

You can't solve this without more information. You can rearrange in terms of k, but you can't solve for k. What is your exact question?

OpenStudy (anonymous):

Change it to 7k=4w+28 then divide by 7

OpenStudy (hlares):

To "solve" for k, you will first want to move the other variable to the right side of the equation. To do that, you will subtract 4w from both sides. After that, you will need to divide both sides by -7. Thus: 4w-7k=28 -7k=28-4w k=(4/7)w-4

OpenStudy (anonymous):

Divide all the numbers by 7

OpenStudy (sawq):

@tom982 The exact question is: Solve 4w − 7k = 28 for k. Show your work.

OpenStudy (sawq):

Thanks guys for all ur help but im getting mixed messages...

OpenStudy (anonymous):

Then it's a badly worded question. @Hlares showed the working out.

OpenStudy (sawq):

why is @meia107 wrong tho?

OpenStudy (sawq):

I just wanna fully understand

OpenStudy (anonymous):

He's not, @Hlares just explained it better :)

OpenStudy (anonymous):

4w-7k=28 -7k=4w+28 ___________ -7 k=-4/7w+-4

OpenStudy (anonymous):

And there's your answer

OpenStudy (anonymous):

:-)

OpenStudy (sawq):

Thanks yall :) I understand now

OpenStudy (anonymous):

Your welcome